OpenStudy (anonymous):

nd the zeros of the function f(x) = 2x3 – 9x + 3

OpenStudy (helder_edwin):

finding the zeros means solving the equation \[ \large f(x)=0 \]

OpenStudy (anonymous):

So: \[2x^3-9x+3=0\] Do you know how to continue?

OpenStudy (anonymous):

a little more please

OpenStudy (anonymous):

I suggest the Newton Method to find the roots.

OpenStudy (anonymous):

\[x_{n+1}=x_n - \frac{f(x)}{f'(x)} \]

OpenStudy (anonymous):

then use iteration until you find a result you are happy with.

OpenStudy (helder_edwin):

@rdiddy212 do u know derivatives to do what @Spacelimbus suggested?

OpenStudy (anonymous):

Refer to attached plot. The roots, calculated by the Mathematica function, NSolve, are the following to 15+ digits:\[\{-2.2716332985527448,0.34224143691160663,1.9293918616411379\} \]